# Comprehensive X-ray Observations of the Exceptional Ultra-long X-ray and Gamma-ray Transient GRB 250702B with Swift, NuSTAR, and Chandra: Insights from the X-ray Afterglow Properties

### v1

@@ 1 Introduction
Gamma-ray bursts (GRBs) are among the most energetic explosions in the Universe, releasing $10^{48-52}$ erg of energy (collimation corrected) in the X-ray and gamma-ray regime (11-$10,000$ keV). They are commonly classified as short or long GRBs based on their prompt emission duration (C. Kouveliotou et al., 1993), and hardness ratio (Nakar2007; O. Bromberg et al., 2013). The prompt duration is typically linked to the lifetime of the central engine after the jet breaks out from the surrounding stellar material (e.g., O. Bromberg et al., 2012, 2013). A subset of GRBs, referred to as ultra-long GRBs, display exceptionally extended prompt gamma-ray emission lasting thousands to tens of thousands of seconds (e.g., B. Gendre et al., 2013; Levan2014; K. Ioka et al., 2016).
@@
Due to its on-sky location’s close proximity to the Galactic plane ($b\sim$55 deg), its nature was initially ambiguous, with possible interpretations ranging from a new X-ray binary in outburst or a peculiar extragalactic transient. While debate over its classification continues (Levan2025; Oganesyan2025; Neights2025; J. Carney et al., 2025; B. P. Gompertz et al., 2025; P. Beniamini et al., 2025; J. Granot, 2025), the most plausible scenarios are that it represents either an ultra-long GRB or a relativistic TDE. The duration of the source’s flaring activity, lasting for at least >>$25$ ks (observer frame) at gamma-ray wavelengths (D. Frederiks et al., 2025; Neights2025), easily places it among the class of ultra-long gamma-ray bursts. However, the Einstein Probe mission identified source activity at soft X-rays (0.50.5-44 keV) starting nearly a full 24 hours before the first gamma-ray trigger (H. Q. Cheng et al., 2025a), which is not easily reconciled with standard GRB models.
Following the arcminute localization by the Einstein Probe (H. Q. Cheng et al., 2025a), X-ray observations with the Swift X-ray Telescope (XRT; D. N. Burrows et al. 2005) identified a rapidly fading X-ray source and provided an arcsecond localization (J. A. Kennea et al., 2025). Near-infrared observations with the Very Large Telescope revealed an extremely red, quickly decaying counterpart (Levan2025). Subsequent Hubble Space Telescope (HST) observations provided conclusive evidence that the transient lies on top of the light of an irregular host galaxy with a negligible probability ($0.1\%$) of chance alignment (Levan2025), confirming the extragalactic nature of GRB 250702B. A spectroscopic redshift of the host galaxy of zz==$1.036$ was obtained by the James Webb Space Telescope (B. P. Gompertz et al., 2025). Notably, the source is spatially offset from the center of the galaxy (Levan2025; J. Carney et al., 2025), complicating a classical TDE interpretation.
We present comprehensive X-ray observations of GRB 250702B with Swift, NuSTAR, and Chandra, aimed at characterizing its temporal and spectral evolution. At first glance, the observed multi-wavelength behavior is consistent with two leading interpretations: i) a relativistic jetted TDE, potentially originating from an offset intermediate-mass black hole (IMBH), or ii) a peculiar ultra-long GRB featuring a soft X-ray precursor. We explore these possibilities in detail, highlighting the conflicting lines of evidence that make this event an outlier that resists a straightforward classification within existing high-energy transient populations.
@@ 2.1.1 Burst Alert Telescope
The Neil Gehrels Swift Observatory (N. Gehrels et al., 2004) Burst Alert Telescope (BAT; S. D. Barthelmy et al., 2005) is a large field of view, ($\sim$22 st) coded mask imager, capable of imaging in the 1414-$195$ keV energy range. Additionally, BAT is sensitive to impulsive emission originating from outside its coded field of view in the 5050-$350$ keV energy range, though at a reduced sensitivity compared to inside the coded field of view.
@@ 2.3 Chandra X-ray Observatory
Another DDT observation (PI: Eyles-Ferris; R. A. J. Eyles-Ferris et al. 2025) occurred starting on 2025-09-05 at 23:26:41 UT, corresponding to 65.5 d after discovery, using ACIS-S for a total of $39.55$ ks. We analyzed this observation in the same manner, finding 12 counts in a $1.5\arcsec$ radius, and obtain an unabsorbed 0.30.3-$10$ keV flux of $(1.15^{+0.40}_{-0.30})\times 10^{-14}$ erg cm-2 s-1.
@@ 4.1 Ultra-long Prompt Emission
GRB 250702B displayed unusual prompt emission properties. For starters, it triggered a variety of high-energy monitors multiple times on July 2, 2025 with a gamma-ray duration in excess of $25$ ks (see Neights2025, for a detailed discussion). Additional high-energy detections were uncovered on July 1, 2025 by the Einstein Probe and Fermi, extending the prompt emission duration to longer than a day (H. Q. Cheng et al., 2025a; Neights2025). In fact, Swift/BAT [^9] As these Swift/BAT detections occur prior to the start of soft X-ray observations with XRT, their interpretation as either additional prompt episodes or X-ray flares is unclear, though see Neights2025.2.1.1; Figure 5) also detected the source in Survey mode on two additional instances that followed the Fermi triggers, further pushing the prompt duration [^10] We note that these Swift/BAT detections are included in the 25 ks determination of the prompt emission duration by Neights2025. (and therefore central engine activity) to extreme values.
As such, the exact start time of the event is unclear, though our analysis favors an afterglow onset timed with the July 2 triggers (Figure 1; right panel). The delay between the July 1 and July 2 detections is unusual for GRBs, which have never before been found to have such long delay between precursor emission and the main burst episode. It is possible that this is a selection bias due to the lack of wide-field soft X-ray monitors like the Einstein Probe in the past. It is therefore possible that some previously known ultra-long GRBs also show such sub-energetic ($10^{48-49}$ erg) precursor emission that went missed.
While these properties are atypical for GRBs, they have likewise never been observed from a relativistic TDE. While relativistic TDEs have been uncovered in binned gamma-ray data (S. B. Cenko et al., 2012; G. C. Brown et al., 2015), or 1000 s long image triggers (Sakamoto2011J1644), they have never triggered high-energy monitors in the standard “rate” trigger method that detects typical GRBs, and they have definitely not done so multiple times in a 24 hour window. As near all-sky gamma-ray monitors have been operating for decades, multiple gamma-ray triggers would be more difficult to miss in the past than a soft X-ray precursor to a GRB that could only have been found by the Einstein Probe which launched only in the last 2 years (Yuan2025). A detailed sub-threshold gamma-ray search in the day(s) preceding ultra-long GRBs is likely necessary to further investigate this possibility.
@@ 4.3 An Internal or External Origin for the X-ray Lightcurve?
The exact cause of the X-ray emission has strong implications for the precise nature of the progenitor system and its central engine. In particular, any central engine capable of reproducing the observed X-ray emission must remain active (for at least) between the initial EP detection (H. Q. Cheng et al., 2025a) and the end of the first NuSTAR observation. In fact, numerical simulations have shown that the soft X-ray emission is a better tracer of the jet activity duration than the gamma-ray duration (Parsotan2024). Cumulatively, the observations require an engine duration in excess of $\gtrsim$33 days in the observer frame, corresponding to $\gtrsim$$1.5$ days ($\gtrsim$$1.6\times 10^{5}$ s) in the rest frame.
@@
While past GRB afterglow observations with NuSTAR do not display this short timescale behavior (§D; Figure 17), this is only a small sample of the GRB phenomena, which is notably diverse. Ultra-long GRBs (e.g., GRB 121027A) do display comparable X-ray variability and dips (Levan2014) out to $\sim$$10^{4}$ d in the rest frame (Figure 10), see, e.g., Figure 14 of P. A. Evans et al. (2014a). This late-time X-ray variability is likely linked to ongoing central engine activity and ongoing prompt gamma-ray emission. However, the timescale required for GRB 250702B is significantly longer than the current ultra-long sample, and has the additional difficulty of producing the intial EP “precursor” emission. We further discuss the prospects for the progenitor of GRB 250702B in §5.
@@ 5.1 An Intermediate-mass Black Hole Progenitor
Associating the observed prompt emission duration (§4.1) with accretion onto the black hole following the tidal disruption of a star, the peak must have occurred on a timescale of at least $\sim$11 d (observer frame), allowing for the initial detection by the Einstein Probe $\sim$$24$ hr before the Fermi triggers (Table 2). If the disrupted star was on the main sequence and completely destroyed, the time to peak emission (from initial pericenter) is $t_{\rm peak}\approx$$25M_{\bullet,6}^{1/2}\textrm{ d }$ (E. R. Coughlin & C. J. Nixon, 2022; A. Bandopadhyay et al., 2024; J. Fancher et al., 2025), where $M_{\bullet,6}$ is the mass of the disrupting black hole in units of $10^{6}M_{\odot}$. After the peak time of the fallback accretion rate (also referred to as the fallback time), the X-ray lightcurve will decay (e.g., $t^{-5/3}$) as observed in past relativistic TDEs (Figure 10). Here, we observe the source to be fading from the initial XRT observation ($\sim$$0.5$ d after the initial triggers). Therefore, we consider a conservative bound on the peak time of fallback accretion to be $t_{\rm peak}\lesssim$$\big(\frac{1}{1+z}\big)1.5$ d. Equating this peak timescale to $\lesssim$$\big(\frac{1}{1+z}\big)1.5$ d then constrains the mass of the disrupting black hole to $\lesssim$$9\times 10^{2}M_{\odot}$, which is firmly in the IMBH regime and could, e.g., be present in a low-mass satellite of the visually distorted massive parent galaxy (see Levan2025; J. Carney et al., 2025). Allowing for an earlier disruption of the star (though disfavored by the existing data) up to a day before even the initial detections (H. Q. Cheng et al., 2025a; Neights2025), would extend this requirement on the black hole mass to $\lesssim$$2.5\times 10^{3}M_{\odot}$. The required black hole masses are consistent with the requirement ($\lesssim$$5\times 10^{4}M_{\odot}$) from the Schwarzchild light crossing time (§4.1). The partial disruption of the star would only lengthen this timescale, and while placing it on a tightly bound orbit could reduce the time to peak for a higher-mass black hole, it would require an additional mechanism, e.g., Hills capture (J. G. Hills, 1988) or disruption by a black hole binary, to generate such an orbit.
@@
Thus, while we can strongly disfavor a WD-IMBH merger, we cannot exclude a MS-IMBH merger (for further discussion, see J. Granot, 2025). The MS-IMBH merger is consistent with many of the observed timescales, but still has difficulties as i) the multiple gamma-ray triggers are unlike any observed in past relativistic TDEs, ii) the X-ray lightcurve is consistent with the luminosities of known GRBs and not known relativistic TDEs (Figures 9 and 10), and iii) the event follows standard GRB prompt-afterglow correlations (Figure 11). Additionally, the observed rate of GRBs, even ultra-long GRBs (though this depends on the exact duration requirement for this definition; see Levan2014), exceeds the known number of relativistic TDEs.
@@ 5.2 A Stellar-mass Black Hole Progenitor
The observed properties of GRB 250702B are most similar to those found in the (notably small) sample of ultra-long GRBs. The main characteristics that make this event appear an outlier are the discovery of un-triggered high-energy emission up to a day prior to its first on-board gamma-ray triggers (H. Q. Cheng et al., 2025a; Neights2025) and the continued short timescale flaring out to 2 days (§4.3). Besides these factors, all other properties appear typical of GRBs. These include: the X-ray luminosity and rapid lightcurve decay, the triggered gamma-ray emission, the isotropic-equivalent gamma-ray energy release, the gamma-ray efficiency, the prompt emission’s minimum variability timescale, the consistency of the multi-wavelength lightcurve with external afterglow emission in a dense wind environment, the dusty host galaxy, and the offset location within its host. In fact, each of these properties are instead atypical for a massive black hole progenitor (§5.1).
@@
However, the discovery of the soft X-ray precursor and the extremely long (>>$25$ ks) gamma-ray emission are difficult to reconcile in a traditional collapsar scenario, even considering a peculiar progenitor such as the collapse of a blue supergiant (Meszaros2001BSG; Stratta2013; Nakauchi2013; K. Ioka et al., 2016). It is worth noting that the Einstein Probe is revealing that soft X-ray emission can extend for 1010-$100$ times the gamma-ray emission for what otherwise would appear a typical GRB at gamma-ray frequencies (see, e.g., Levan2024; Yin2024; Liu2024; Yin2025ep250404a). If this same phenomena can be extended to ultra-long GRBs, it could potentially explain the longer duration of the prompt emission. Previously existing selection biases, based on the lack of wide-field soft X-ray telescopes, may have played a role in limiting the discovery of soft precursor emission. It is possible that the initial soft X-ray emission detected by the Einstein Probe (H. Q. Cheng et al., 2025a) is due to the formation of a cocoon (as in, e.g., Leung2025) as the jet (or multiple shells launched by the central engine) attempts to breakout of the extended stellar radius of a blue supergiant. As subsequent shells launched within the jet do not have as great a difficulty burrowing out of the stellar material, we have a clearer picture of their prompt gamma-rays (likely produced through internal shocks). However, the requirements for the duration of the central engine are still extreme even in such a scenario, and lead us to conclude that a collapsar (even from a blue supergiant) is an unlikely progenitor for GRB 250702B.
@@
Based on all of the considerations presented in this manuscript, we find that while GRB 250702B displays similar properties of both classical and ultra-long GRBs (e.g., Figure 11), with the main exception being the early pre-trigger “precursor” emission, it may be explained by a “hybrid” progenitor system involving a stellar-mass black hole, but with a TDE-like central engine. This is also favored by the consistency of the surrounding environment with the profile of a stellar wind. The main possibilities for this hybrid system are the merger of a helium star with a stellar-mass black hole (C. L. Fryer & S. E. Woosley, 1998; Woosley2012; Neights2025) or a micro-TDE by a stellar-mass black hole (Perets2016; P. Beniamini et al., 2025). These scenarios have been shown to be compatible with day long central engine activity as required by this event and may more naturally explain the precursor emission, 1 s minimum variability timescale of the prompt emission, long-lived central engine ($\gtrsim$33 days), and short timescale X-ray flaring produced by an accretion driven relativistic outflow. These progenitor scenarios can also potentially explain the current sample of ultra-long GRBs. A complete discussion of the micro-TDE scenario is provided by P. Beniamini et al. (2025), and the helium merger scenario is discussed by Neights2025.

### v2

@@ 1 Introduction
Gamma-ray bursts (GRBs) are among the most energetic explosions in the Universe, releasing $10^{48-52}$ erg of energy (collimation corrected) in the X-ray and gamma-ray regime (11-$10,000$ keV). They are commonly classified as short or long GRBs based on their prompt emission duration (C. Kouveliotou et al., 1993), and hardness ratio (Nakar2007; O. Bromberg et al., 2013). The prompt duration is typically linked to the lifetime of the central engine after the jet breaks out from the surrounding stellar material (e.g., O. Bromberg et al., 2012, 2013). A subset of GRBs, referred to as ultra-long GRBs, display exceptionally extended prompt gamma-ray emission lasting thousands to tens of thousands of seconds (e.g., 2005AstL...31..291T; B. Gendre et al., 2013; Levan2014; K. Ioka et al., 2016).
@@
Due to its on-sky location’s close proximity to the Galactic plane ($b\sim$55 deg), its nature was initially ambiguous, with possible interpretations ranging from a new X-ray binary in outburst or a peculiar extragalactic transient. While debate over its classification continues (Levan2025; Oganesyan2025; Neights2025; J. Carney et al., 2025; B. P. Gompertz et al., 2025; P. Beniamini et al., 2025; J. Granot et al., 2025), the most plausible scenarios are that it represents either an ultra-long GRB or a relativistic TDE. The duration of the source’s flaring activity, lasting for at least >>$25$ ks (observer frame) at gamma-ray wavelengths (D. Frederiks et al., 2025; Neights2025), easily places it among the class of ultra-long gamma-ray bursts. However, the Einstein Probe mission identified source activity at soft X-rays (0.50.5-44 keV) starting nearly a full 24 hours before the first gamma-ray trigger (H. Q. Cheng et al., 2025a; EP250702a-arxiv), which is not easily reconciled with standard GRB models.
Following the arcminute localization by the Einstein Probe (H. Q. Cheng et al., 2025a; EP250702a-arxiv), X-ray observations with the Swift X-ray Telescope (XRT; D. N. Burrows et al. 2005) identified a rapidly fading X-ray source and provided an arcsecond localization (J. A. Kennea et al., 2025). Near-infrared observations with the Very Large Telescope revealed an extremely red, quickly decaying counterpart (Levan2025). Subsequent Hubble Space Telescope (HST) observations provided conclusive evidence that the transient lies on top of the light of an irregular host galaxy with a negligible probability ($0.1\%$) of chance alignment (Levan2025), confirming the extragalactic nature of GRB 250702B. A spectroscopic redshift of the host galaxy of zz==$1.036$ was obtained by the James Webb Space Telescope (B. P. Gompertz et al., 2025). Notably, the source is spatially offset from the center of the galaxy (Levan2025; J. Carney et al., 2025), complicating a classical TDE interpretation.
We present comprehensive X-ray observations of GRB 250702B with Swift, NuSTAR, and Chandra, aimed at characterizing its temporal and spectral evolution. At first glance, the observed multi-wavelength behavior is consistent with two leading interpretations: i) a relativistic jetted TDE, potentially originating from an offset intermediate-mass black hole (IMBH), or ii) a peculiar ultra-long GRB featuring a soft X-ray “precursor”. We explore these possibilities in detail, highlighting the conflicting lines of evidence that make this event an outlier that resists a straightforward classification within existing high-energy transient populations.
@@ 2.1.1 Burst Alert Telescope
The Neil Gehrels Swift Observatory (N. Gehrels et al., 2004) Burst Alert Telescope (BAT; S. D. Barthelmy et al., 2005) is a large field of view, ($\sim$22 sr) coded mask imager, capable of imaging in the 1414-$195$ keV energy range. Additionally, BAT is sensitive to impulsive emission originating from outside its coded field of view in the 5050-$350$ keV energy range, though at a reduced sensitivity compared to inside the coded field of view.
@@ 2.3 Chandra X-ray Observatory
Another DDT observation (PI: Eyles-Ferris; R. A. J. Eyles-Ferris et al. 2025; R. A. Eyles-Ferris et al. 2025) occurred starting on 2025-09-05 at 23:26:41 UT, corresponding to 65.5 d after discovery, using ACIS-S for a total of $39.55$ ks. We analyzed this observation in the same manner, finding 12 counts in a $1.5\arcsec$ radius, and obtain an unabsorbed 0.30.3-$10$ keV flux of $(1.15^{+0.40}_{-0.30})\times 10^{-14}$ erg cm-2 s-1.
@@ 4.1 Ultra-long Prompt Emission
GRB 250702B displayed unusual prompt emission properties. For starters, it triggered a variety of high-energy monitors multiple times on July 2, 2025 with a gamma-ray duration in excess of $25$ ks (see Neights2025, for a detailed discussion). Additional high-energy detections were uncovered on July 1, 2025 by the Einstein Probe and Fermi, extending the prompt emission duration to longer than a day (H. Q. Cheng et al., 2025a; EP250702a-arxiv; Neights2025). In fact, Swift/BAT [^9] As these Swift/BAT detections occur prior to the start of soft X-ray observations with XRT, their interpretation as either additional prompt episodes or X-ray flares is unclear, though see Neights2025.2.1.1; Figure 5) also detected the source in Survey mode on two additional instances that followed the Fermi triggers, further pushing the prompt duration [^10] We note that these Swift/BAT detections are included in the 25 ks determination of the prompt emission duration by Neights2025. (and therefore central engine activity) to extreme values.
As such, the exact start time of the event is unclear, though our analysis favors an afterglow onset timed with the July 2 triggers (Figure 1; right panel). The delay between the July 1 and July 2 detections is unusual for GRBs, which have never before been found to have such long delay between the early “precursor” emission and the main burst episode. It is possible that this is a selection bias due to the lack of wide-field soft X-ray monitors like the Einstein Probe in the past. It is therefore possible that some previously known ultra-long GRBs also show such sub-energetic precursor emission that went missed. However, the Einstein Probe lightcurve shows a smooth rise to peak emission (EP250702a-arxiv), coinciding with the final Fermi trigger (Neights2025), which excludes its interpretation as a “precursor”. In any case, such underlying soft X-ray emission would easily be missed in previous ultra-long events.
While these properties are atypical for GRBs, they have likewise never been observed from a relativistic TDE. While relativistic TDEs have been uncovered in binned gamma-ray data (S. B. Cenko et al., 2012; G. C. Brown et al., 2015), or 1000 s long image triggers (Sakamoto2011J1644), they have never triggered high-energy monitors in the standard “rate” trigger method that detects typical GRBs, and they have definitely not done so multiple times in a 24 hour window. As near all-sky gamma-ray monitors have been operating for decades, multiple gamma-ray triggers would be more difficult to miss in the past than early soft X-ray emission that could only have been found by the Einstein Probe which launched only in the last 2 years (Yuan2025). A detailed sub-threshold gamma-ray search in the day(s) preceding ultra-long GRBs is likely necessary to further investigate this possibility.
@@ 4.3 An Internal or External Origin for the X-ray Lightcurve?
The exact cause of the X-ray emission has strong implications for the precise nature of the progenitor system and its central engine. In particular, any central engine capable of reproducing the observed X-ray emission must remain active (for at least) between the initial EP detection (H. Q. Cheng et al., 2025a; EP250702a-arxiv) and the end of the first NuSTAR observation. In fact, numerical simulations have shown that the soft X-ray emission is a better tracer of the jet activity duration than the gamma-ray duration (Parsotan2024). Cumulatively, the observations require an engine duration in excess of $\gtrsim$33 days in the observer frame, corresponding to $\gtrsim$$1.5$ days ($\gtrsim$$1.6\times 10^{5}$ s) in the rest frame.
@@
While past GRB afterglow observations with NuSTAR do not display this short timescale behavior (§D; Figure 17), this is only a small sample of the GRB phenomena, which is notably diverse. Ultra-long GRBs (e.g., GRB 121027A) do display comparable X-ray variability and dips (Levan2014) out to $\sim$$10^{4}$ s in the rest frame (Figure 10), see, e.g., Figure 14 of P. A. Evans et al. (2014a). This late-time X-ray variability is likely linked to ongoing central engine activity and ongoing prompt gamma-ray emission. However, the timescale required for GRB 250702B is significantly longer than the current ultra-long sample, and has the additional difficulty of producing the initial EP emission (EP250702a-arxiv). We further discuss the prospects for the progenitor of GRB 250702B in §5.
@@ 5.1 An Intermediate-mass Black Hole Progenitor
Associating the observed prompt emission duration (§4.1) with accretion onto the black hole following the tidal disruption of a star, the peak must have occurred on a timescale of at least $\sim$11 d (observer frame), allowing for the initial detection by the Einstein Probe $\sim$$24$ hr before the Fermi triggers (Table 2). If the disrupted star was on the main sequence and completely destroyed, the time to peak emission (from initial pericenter) is $t_{\rm peak}\approx$$25M_{\bullet,6}^{1/2}\textrm{ d }$ (E. R. Coughlin & C. J. Nixon, 2022; A. Bandopadhyay et al., 2024; J. Fancher et al., 2025), where $M_{\bullet,6}$ is the mass of the disrupting black hole in units of $10^{6}M_{\odot}$. After the peak time of the fallback accretion rate (also referred to as the fallback time), the X-ray lightcurve will decay (e.g., $t^{-5/3}$) as observed in past relativistic TDEs (Figure 10). Here, we observe the source to be fading from the initial XRT observation ($\sim$$0.5$ d after the initial triggers). Therefore, we consider a conservative bound on the peak time of fallback accretion to be $t_{\rm peak}\lesssim$$\big(\frac{1}{1+z}\big)1.5$ d. Equating this peak timescale to $\lesssim$$\big(\frac{1}{1+z}\big)1.5$ d then constrains the mass of the disrupting black hole to $\lesssim$$9\times 10^{2}M_{\odot}$, which is firmly in the IMBH regime and could, e.g., be present in a low-mass satellite of the visually distorted massive parent galaxy (see Levan2025; J. Carney et al., 2025). Allowing for an earlier disruption of the star (though disfavored by the existing data) up to a day before even the initial detections (H. Q. Cheng et al., 2025a; EP250702a-arxiv; Neights2025), would extend this requirement on the black hole mass to $\lesssim$$2.5\times 10^{3}M_{\odot}$. The required black hole masses are consistent with the requirement ($\lesssim$$5\times 10^{4}M_{\odot}$) from the Schwarzchild light crossing time (§4.1). The partial disruption of the star would only lengthen this timescale, and while placing it on a tightly bound orbit could reduce the time to peak for a higher-mass black hole, it would require an additional mechanism, e.g., Hills capture (J. G. Hills, 1988) or disruption by a black hole binary, to generate such an orbit.
@@
Thus, while we can strongly disfavor a WD-IMBH merger, we cannot exclude a MS-IMBH merger (for further discussion, see J. Granot et al., 2025). The MS-IMBH merger is consistent with many of the observed timescales, but still has difficulties as i) the multiple gamma-ray triggers are unlike any observed in past relativistic TDEs, ii) the X-ray lightcurve is consistent with the luminosities of known GRBs and not known relativistic TDEs (Figures 9 and 10), and iii) the event follows standard GRB prompt-afterglow correlations (Figure 11). Additionally, the observed rate of GRBs, even ultra-long GRBs (though this depends on the exact duration requirement for this definition; see Levan2014), exceeds the known number of relativistic TDEs.
@@ 5.2 A Stellar-mass Black Hole Progenitor
The observed properties of GRB 250702B are most similar to those found in the (notably small) sample of ultra-long GRBs. The main characteristics that make this event appear an outlier are the discovery of un-triggered high-energy emission up to a day prior to its first on-board gamma-ray triggers (H. Q. Cheng et al., 2025a; EP250702a-arxiv; Neights2025) and the continued short timescale flaring out to 2 days (§4.3). Besides these factors, all other properties appear typical of GRBs. These include: the X-ray luminosity and rapid lightcurve decay, the triggered gamma-ray emission, the isotropic-equivalent gamma-ray energy release, the gamma-ray efficiency, the prompt emission’s minimum variability timescale, the consistency of the multi-wavelength lightcurve with external afterglow emission in a dense wind environment, the dusty host galaxy, and the offset location within its host. In fact, each of these properties are instead atypical for a massive black hole progenitor (§5.1).
@@
However, the discovery of the early soft X-ray emission (EP250702a-arxiv) and the extremely long (>>$25$ ks; Neights2025) gamma-ray emission are difficult to reconcile in a traditional collapsar scenario, even considering a peculiar progenitor such as the collapse of a blue supergiant (Meszaros2001BSG; Stratta2013; Nakauchi2013; K. Ioka et al., 2016). It is worth noting that the Einstein Probe is revealing that soft X-ray emission can extend for 1010-$100$ times the gamma-ray emission for what otherwise would appear a typical GRB at gamma-ray frequencies (see, e.g., Levan2024; Yin2024; Liu2024; Yin2025ep250404a). If this same phenomena can be extended to ultra-long GRBs, it could potentially explain the longer duration of the prompt emission. Previously existing selection biases, based on the lack of wide-field soft X-ray telescopes, may have played a role in limiting the discovery of soft precursor emission. It is possible that the initial soft X-ray emission detected by the Einstein Probe (H. Q. Cheng et al., 2025a; EP250702a-arxiv) is due to the formation of a cocoon (as in, e.g., Leung2025) as the jet (or multiple shells launched by the central engine) attempts to breakout of the extended stellar radius of a blue supergiant. As subsequent shells launched within the jet do not have as great a difficulty burrowing out of the stellar material, we have a clearer picture of their prompt gamma-rays (likely produced through internal shocks). However, the requirements for the duration of the central engine are still extreme even in such a scenario, and lead us to conclude that a collapsar (even from a blue supergiant) is an unlikely progenitor for GRB 250702B.
@@
Based on all of the considerations presented in this manuscript, we find that while GRB 250702B displays similar properties of both classical and ultra-long GRBs (e.g., Figure 11), with the main exception being the early pre-trigger X-ray emission (EP250702a-arxiv), it may be explained by a “hybrid” progenitor system involving a stellar-mass black hole, but with a TDE-like central engine. This is also favored by the consistency of the surrounding environment with the profile of a stellar wind. The main possibilities for this hybrid system are the merger of a helium star with a stellar-mass black hole (C. L. Fryer & S. E. Woosley, 1998; Woosley2012; Neights2025) or a micro-TDE by a stellar-mass black hole (Perets2016; P. Beniamini et al., 2025). These scenarios have been shown to be compatible with day long central engine activity as required by this event and may more naturally explain the early X-ray emission, 1 s minimum variability timescale of the prompt emission, long-lived central engine ($\gtrsim$33 days), and short timescale X-ray flaring produced by an accretion driven relativistic outflow. These progenitor scenarios can also potentially explain the current sample of ultra-long GRBs. A complete discussion of the micro-TDE scenario is provided by P. Beniamini et al. (2025), and the helium merger scenario is discussed by Neights2025.